A page for a journal contains 70 square inches of type written words. The height of a page is twice the width. If the Margin around the type is to be 2 inches uniformly, what are the dimensions of a page?
Answer by mananth(16946)

area of written area = 70 in^2
let width be x
height = 2x
2x^2= 70
x^2=35
x=sqrt35= 5.91"
...
margin all round = 2 "
width of page = 5.91"+4 " = 9.91
height of page = 2*(5.91)+4 "= 15.8
width of page = 9.91"
height = 15.8"